how do i solve in detail this problem:
3x^2-8x=16

Answer by checkley77(12844)

3x^2-8x=16
3x^2-8x-16=0
(3x+4)(x-4)=0
3x+4=0
3x=-4
x=-4/3 ans.
x-4=0
x=4 ans.